OpenStudy (anonymous):

at an accident scene, a police officer may determine the rate of speed, r, in miles per hour of the car by using the following fromula r= radical symbol 20l, where l is the length of the skid marks. How fast was a car going if the skid marks at the scene are 180 feet long?

OpenStudy (chaise):

\[r = \sqrt{20L}\] r = sqrt(20x180) r = sqrt(3600) r = 60 mph

OpenStudy (anonymous):

thanks